Evaluate (-2x)2 and -2x 2 when x = 3
Both of the second 2's in each problem is one of those little numbers which i believe mean to the 2nd power??

Answer by nycsharkman(136)

Evaluate (-2x)^2 and -2x^2 when x = 3.
If x = 3, then replace x with 3 and simplify.
(-2(3))^2 = (-6)^2 = 36
===========================================
For the next question, notice that there is no parentheses. This means that only the x value is squared and then the product is multiplied by -2.
-2(3)^2 = -2(9) = -18
Is this clear?